OpenStudy (anonymous):

If f'(x)=2/x and f(sqrt(e))=5, then f(e)=

OpenStudy (anonymous):

\[ dy={2\over x}dx\implies y=2\ln(x)+C=f(x)\\ \] to get "C", use the condition \[f(\sqrt{e})=2\ln(\sqrt{e})+C=5\\\implies{\cancel{2}\over\cancel{2}}\ln(e)+C=5\implies 1+C=5\implies C=4\\ \therefore f(x)=2\ln(x)+4 \] now substitute x=e and evaluate!

OpenStudy (anonymous):

Got it.